# Thermally-induced mimicry of quantum cluster excitations and implications for the magnetic transition in FePX3

### v1

@@ Abstract
In two dimensional magnets, the interplay of thermal fluctuations and spin anisotropy control the existence of long-range magnetic order. In the van der Waals antiferromagnets FePX3, orbital degeneracy in the $t_{2g}$ levels of the Fe2+ ions in octahedral coordination yields strong uniaxial anisotropy, which stabilizes magnetic order up to $T\approx 100$ K. Recent inelastic neutron scattering measurements around the magnetic ordering transition have shown the existence of a broad spectrum of magnetic fluctuations with nontrivial momentum dependence, what has been interpreted as evidence for localized entangled cluster excitations. In this paper, we offer an alternative interpretation using classical non-linear spin dynamics simulations. We present stochastic Landau Lifshitz dynamics simulations that reproduce the neutron scattering measurements of Chen et al. [1] on FePX3. These calculations faithfully explain the dynamical structure factor’s momentum and energy dependence and point to a classical origin for the excitations observed in neutron spectroscopy and that the order-disorder transition can be understood in terms of thermal fluctuations overcoming the anisotropy energy.
@@ I Introduction
In this paper, we discuss the nature of the spin fluctuations in the Fe2+ compounds, (FePSe3 and FePS3), with specific emphasis on the former compound in light of recent comprehensive experimental results [1]. It has been suggested that the spin excitations in the correlated paramagnetic regime are consistent with transitions between quantum-correlated hexagonal spin clusters, as discussed in the context of large-$S$ pyrochlore Heisenberg antiferromagnets [38, 39, 40, 41]. Here, we detail an alternative explanation as classical spin dynamics exhibiting non-linearities from thermal fluctuations. This interpretation overcomes the explanatory difficulty that thermal fluctuations are antithetical to forming spatially extended entangled clusters. Moreover, our finite temperature calculations shed further light on the nature of the order-disorder phase transition in the FePX3 family of materials.
@@ II Exchange Model
The measured magnetic moment for both compounds is close to the spin-only value, suggesting that the strong trigonal distortion largely quenches the orbital moment; this motivates a traditional treatment of the local spin space as purely dipolar (using coherent states of $SU(2)$) in the absence of off-diagonal magnetic exchange interactions, with $S=2$. Still, small discrepancies between the experimentally measured spectra and predictions from theoretical spin models remain. Recently, in the case of FePS3, it has been suggested that these discrepancies arise due to a partial mixing of excited $J=2$ and $J=3$ spin-orbital levels since an orbital degeneracy exists in the $t_{2g}$ levels of Fe2+ [42, 43].
We first discuss the exchange model of FePSe3 before commenting on to FePS3. In FePSe3, the magnetic ordering wavevector $\boldsymbol{k}=(\frac{1}{2},0,\frac{1}{2})$ describes a zig-zag order in the plane and antiferromagnetic out of plane order. As shown in Ref. 1, the out of plane dispersion is significantly weaker than the in-plane dispersion, we therefore set the nearest-neighbor out of plane coupling $J_{c}$ equal to a small value ($J_{c}=0.01$ meV) to achieve the correct ordering wavevector. We consider the two models of Ref. 1 consistent with the crystal symmetry to describe the in-plane magnetic interactions. The first, labeled BL, is a $J_{1}$-$J_{2}$-$J_{3}$ model with single-ion anisotropy, $D$. The second, labeled BQ, is a $J_{1}$-$J_{2}$-$J_{3}$-$K_{1}$ model, where $K_{1}$ is a nearest neighbor ferromagnetic biquadratic interaction. This last term introduces inequivalence in the exchange between different Bravais lattices [31] and may originate from magnetoelastic coupling or from the presence of a finite orbital moment leading to non-zero multipolar exchange as discussed in Ref. 44 and recently in Ref.42 in the context of FePS3. The Hamiltonian can thus be written as
@@
In the case of FePS3, a $J_{1}$-$J_{2}$-$J_{3}$-$K_{1}$-$K_{2}$ model with single-ion anisotropy best describes the inelastic neutron scattering data [31], with parameters qualitatively similar to the fitted values for FePSe3 [1] including with a sizable ferromagnetic nearest-neighbor and antiferromagnetic next-nearest-neighbor Heisenberg exchange, a strong uniaxial anisotropy, and a ferromagnetic biquadratic nearest-neighbor exchange.
As similar Hamiltonians are reported for the two FePX3 systems, we only present calculations for FePSe3 in this work, but for completeness, we list the proposed parameters for both compounds in Table 1. Subtle differences may arise between the two compounds due to the different atomic numbers of the chalcogen ions, which affects both the strength of spin-orbit coupling and orbital overlap.
Throughout this paper, we utilize the Renormalized Classical Spin (RCS) theory [45] to describe the spin dynamics. This approach more accurately describes single-ion anisotropies and biquadratic exchange than the traditional large-$S$ theory, because it directly uses the classical limit of higher-order multipolar spin operators, rather than taking the classical limit of each individual spin operator. For quadratic anisotropies, the correspondence between the two methods is exact in the limit $S\to\infty$, but a correction factor must be applied for finite $S$. In the case of biquadratic exchange, the RCS theory generates a correction that is bilinear in spin operators and must be accounted for in the strength of exchange interactions to match the fitted spectrum in Ref 1. This term is independent of $S$ and so even for small values of $1/S$, this correction to the large-$S$ limit remains. The correspondence between large-$S$ and RCS is summarized in Appendix A. These renormalization factors are distinct from the quantum corrections generated by a mean-field averaging of higher-order $1/S$ corrections to the Hamiltonian, which have been neglected from our analysis.
@@ III Static Correlations
We start by determining the ground state spin configuration of FePSe3 by using the (adequately classically renormalized) exchanges listed in Table 1. We do so unbiasedly by performing Langevin dynamics simulations using the Sunny.jl package [46, 47, 48]. Starting from the high temperature (random) limit, we thermalize a (24,24,2) supercell of spins (in the hexagonal setting) followed by gradient descent to find the $T=0$ magnetic structure. We verify that for both FePSe3 models (BL and BQ), the ground state magnetic order is the $\boldsymbol{k}=(\frac{1}{2},0,\frac{1}{2})$ zig-zag structure, or its domains generated under $120^{\circ}$ rotations. In the case of FePS3, we find the $\boldsymbol{k}=(0,1,\frac{1}{2})$ structure, which is the same zig-zag arrangement as observed in FePSe3 up to a redefinition of the basis vectors as appropriate for the $C2/m$ monoclinic space group.
@@
To further explore the nature of the magnetic ordering transition in FePSe3 we calculate the magnetic order parameter as a function of temperature using the same Langevin dynamics approach. Care was taken to ensure samples were taken after sufficient time to thermalize and decorrelate the spin configurations [Appendix  B]. We observe a sharp transition at $T_{N}$ followed by a gradual saturation as $T\to 0$ [Appendix C].
@@
A well-known deficiency of classical simulations is the failure to capture the first-order nature of the transition. This fact was pointed out in Ref. 49 and can be overcome by restricting the dimensionality of the fluctuations manifold by considering Ising spins (and thus taking $J_{i}^{\alpha\beta}\to 0$ for $\alpha,\beta\neq z$). This leads to a better accounting of the first-order nature of this transition. However, such a model does not accurately describe the dynamics in systems with Heisenberg exchange.
@@
Figure 1: The magnetic order parameter corresponding to $\boldsymbol{k}=(\frac{1}{2},0,\frac{1}{2}$) for the biquadratic (BQ) model of FePSe3 simulated using Langevin dynamics on a (24,24,2) supercell. A sharp phase transition is observed at $T_{N}^{\mathrm{cl}}=70$ K. The in-plane magnetic structure is displayed in the inset with up (down) spins in purple (yellow). The first three nearest neighbor bonds, $J_{1}$ (black), $J_{2}$ (red) and $J_{3}$ (blue) are also plotted.
The overestimation of fluctuations close to the ordering temperature also manifests in an underestimation of $T_{\rm N}$. In our classical simulations, we find that $T^{\mathrm{cl}}_{N}=70$ K, whilst experimentally, the ordering temperature was determined to be $T^{\mathrm{exp}}_{\rm N}=110$ K. Notwithstanding these limitations, our classical simulations correctly capture the sharp order parameter increase between a paramagnetic state and a zig-zag antiferromagnetic state below $T_{\rm N}$.
@@ IV Linear Spin Wave Spectrum
A complete understanding of the spin dynamics of FePSe3 requires modeling dispersive higher energy spin-orbital excitations observed around $\sim$ 39 and 44 meV. In principle, this can be achieved by representing the local Hilbert space using coherent states of $SU(5)$ and including biquadratic exchange interactions that directly couple magnetic multipoles. However, the coupling between high- and low-energy excitations is small, leading to a negligible neutron scattering intensity for these modes. Including mixing between higher spin-orbital states could potentially give rise to non-zero matrix elements between the ground-state and higher energy excited states; neutron scattering measurements at higher energy transfer may shed light on this question.
@@
Table 1: Parameters used in our simulations, adapted using the RCS to be consistent with Chen et al. [1] and Wildes et al. [31]. A small interchain coupling has been added to reproduce the reported magnetic ground state. ∗∗Reported value includes the RCS renormalization. Exchange couplings from Ref. 31 include a factor of two to account for the difference in Hamiltonian definition.
FePSe3 FePS3
Bond BL BQ BQ
$J_{1}$2.40 -1.32 -2.01750
$J_{2}$0.23  0.12  0.28875
$J_{3}$  2.01  1.28  0.90000
$J_{c}$  0.01  0.01  0.01460
$K_{1}$  0.00 -0.22∗∗ -0.20400
$K_{2}$  0.00  0.00 -0.03800
$D$ 3.65  3.08  3.36000
@@ V Temperature-dependent spin fluctuations
With this, we seek to address the nature of the spin fluctuations near the magnetic phase transition by simulating the energy-resolved inelastic response as a function of temperature. In Ref 1, neutron spectra were measured at temperatures above and below the phase transition: $T=0.92T_{N}$, $0.97T_{N}$, $1.03T_{N}$ and $1.14T_{N}$. The relevant energy scale that governs the short-range correlations around $T_{N}$ is the magnitude of the order parameter, which acts as a temperature-dependent molecular mean field on dipoles in the presence of thermal fluctuations. The sharp nature of the first-order phase transition and uncertainty associated with the sample temperature control makes it difficult to quantitatively extract the molecular mean field associated with each given temperature. We, therefore, treat the temperature as an effectively free parameter that we adjust to match the excitation spectrum. For instance, well below $T=T_{N}$, the classical simulation over-estimates the reduction of the order parameter due to thermal fluctuations [Fig. 1] and so to reproduce the experimental data we reduce the temperature below the values used in experiment.
@@
Figure 3: (a-g) Simulated inelastic neutron scattering intensity represented as $(\mathbf{Q},E)$ slices taken at increasing temperatures. Slices in the second row are plotted on a smaller intensity scale to aid visibility. Gray vertical lines indicate the location of the M points in the first Brillouin zone. Numerical divergences that appear as artifacts in the ordered phase due to the finite length of the real-time trajectories are deliberately obscured by these gray lines. (g) Constant energy cuts at $E=5$ meV taken through the $(\mathbf{Q},E)$ slice plotted in (a-f).
@@
Simulation results are plotted in Fig. 3. At $T=0.64T_{N}^{\mathrm{cl}}$, a dispersive mode is still visible, albeit with a broadened lineshape compared to the zero-temperature limit. The broad linewidth indicates a finite lifetime induced by thermal nonlinearities of the classical dynamics. At $T=0.86T_{N}^{\mathrm{cl}}$ the dispersive modes are still present with a reduction in intensity reflecting a redistribution of spectral weight to diffuse quasi-elastic features. A further reduction of the anisotropy gap is caused by a thermal fluctuation induced reduction of the ordered moment. As $T$ crosses the ordering temperature [Fig. 3d], broad features extend from the wavevectors associated with the antiferromagnetic ordering wavevector, indicating the presence of short-range antiferromagnetic correlations. These extend up and join the spin wave branch at $T_{N}$. The broad columnar features persist to higher temperatures [Figs. 3d-f] with the increased temperature redistributing intensity to the columns protruding from the locations of the magnetic Bragg peaks. The scattering remains structured at temperatures far exceeding the magnetic ordering temperature [Fig.3f], reproducing the signal observed experimentally [1]. To further clarify this temperature evolution, we calculate a constant energy cut at $E\!=\!5$ meV. At $T=0.64T_{N}^{\mathrm{cl}}$ no signal is observed in this range. As temperature is increased intensity accumulates at the antiferromagnetic wavevector leading to broad features across the Brilloin zone at $T=1.07T_{N}^{\mathrm{cl}}$, $T=1.08T_{N}^{\mathrm{cl}}$ and $T=1.14T_{N}^{\mathrm{cl}}$.

### v3

# Thermally-induced mimicry of quantum cluster excitations and implications for the magnetic transition in FePSe3

@@ Abstract
In two dimensional magnets, the interplay of thermal fluctuations and spin anisotropy control the existence of long-range magnetic order. In the van der Waals antiferromagnets FePX3, orbital degeneracy in the $t_{2g}$ levels of the Fe2+ ions in octahedral coordination yields strong uniaxial anisotropy, which stabilizes magnetic order up to $T\approx 100$ K. Recent inelastic neutron scattering measurements around the magnetic ordering transition have shown the existence of a broad spectrum of magnetic fluctuations with nontrivial momentum dependence, what has been interpreted as evidence for localized entangled cluster excitations. In this paper, we offer an alternative interpretation using classical non-linear spin dynamics simulations. We present stochastic Landau Lifshitz dynamics simulations that reproduce the neutron scattering measurements of Chen et al. [1] on FePSe3. These calculations faithfully explain the dynamical structure factor’s momentum and energy dependence and point to a classical origin for the excitations observed in neutron spectroscopy and that the order-disorder transition can be understood in terms of thermal fluctuations overcoming the anisotropy energy.
@@ I Introduction
In this paper, we discuss the nature of the spin fluctuations in the Fe2+ compounds, (FePSe3 and FePS3), with specific emphasis on the former compound in light of recent comprehensive experimental results [1] [we discuss these results in the context of available measurements of FePS3 in Appendix A]. It has been suggested that the spin excitations in the correlated paramagnetic regime are consistent with transitions between quantum-correlated hexagonal spin clusters, as discussed in the context of large-$S$ pyrochlore Heisenberg antiferromagnets [38, 39, 40, 41]. Here, we detail an alternative explanation as classical spin dynamics exhibiting non-linearities from thermal fluctuations. This interpretation overcomes the explanatory difficulty that thermal fluctuations are antithetical to forming spatially extended entangled clusters. Moreover, our finite temperature calculations shed further light on the nature of the order-disorder phase transition in FePSe3.
@@ II Exchange Model
The measured magnetic moment for FePSe3 is close to the spin-only value, suggesting that the strong trigonal distortion largely quenches the orbital moment; this motivates a traditional treatment of the local spin space as purely dipolar (using coherent states of SU$(2)$) in the absence of off-diagonal magnetic exchange interactions, with $S=2$. Still, small discrepancies between the experimentally measured spectra and predictions from theoretical spin models remain. Recently, in the case of FePS3, it has been suggested that these discrepancies arise due to a partial mixing of excited $J=2$ and $J=3$ spin-orbital levels since an orbital degeneracy exists in the $t_{2g}$ levels of Fe2+ [42, 43].
We now discuss the exchange model of FePSe3 [further comments on FePS3 can be found in Appendix A]. In FePSe3, the magnetic ordering wavevector $\boldsymbol{k}=(\frac{1}{2},0,\frac{1}{2})$ describes a zig-zag order in the plane and antiferromagnetic out of plane order. As shown in Ref. 1, the out of plane dispersion is significantly weaker than the in-plane dispersion, we therefore set the nearest-neighbor out of plane coupling $J_{c}$ equal to a small value ($J_{c}=0.01$ meV) to achieve the correct ordering wavevector. We consider the two models of Ref. 1 consistent with the crystal symmetry to describe the in-plane magnetic interactions. The first, labeled BL (bilinear), is a $J_{1}$-$J_{2}$-$J_{3}$ model with single-ion anisotropy, $D$. The second, labeled BQ (biquadratic), is a $J_{1}$-$J_{2}$-$J_{3}$-$K_{1}$ model, where $K_{1}$ is a nearest neighbor ferromagnetic biquadratic interaction. This last term introduces inequivalence in the exchange between different Bravais lattices [31] and may originate from magnetoelastic coupling or from the presence of a finite orbital moment leading to non-zero multipolar exchange as discussed in Ref. 44 and recently in Ref.42 in the context of FePS3. The Hamiltonian can thus be written as
@@
Similar Hamiltonians are reported for the two FePX3 systems. We only present calculations for FePSe3 in this work, but for completeness, we list the proposed parameters for FePS3 Table 1 based on Refs. [31, 1]. Whilst the low temperature Hamiltonian parameters are similar for both compounds, subtle differences may arise between the two compounds due to the different atomic numbers of the chalcogen ions, which affects both the strength of spin-orbit coupling and orbital overlap. Furthermore, differences in the magnetoelastic properties may play a role close to the ordering transition which is accompanied by an observable structural transition in FePS3 [45].
Throughout this paper, we utilize the Renormalized Classical Spin (RCS) theory [46] to describe the spin dynamics. This approach more accurately describes single-ion anisotropies and biquadratic exchange than the traditional large-$S$ theory, because it directly uses the classical limit of higher-order multipolar spin operators, rather than taking the classical limit of each individual spin operator. For quadratic anisotropies, the correspondence between the two methods is exact in the limit $S\to\infty$, but a correction factor must be applied for finite $S$. In the case of biquadratic exchange, the RCS theory generates a correction that is bilinear in spin operators and must be accounted for in the strength of exchange interactions to match the fitted spectrum in Ref 1. This term is independent of $S$ and so even for small values of $1/S$, this correction to the large-$S$ limit remains. The correspondence between large-$S$ and RCS is summarized in Appendix B. These renormalization factors are distinct from the quantum corrections generated by a mean-field averaging of higher-order $1/S$ corrections to the Hamiltonian, which have been neglected from our analysis. The origin of the latter corrections are a result of decoupling of terms of magnon scattering terms [47].
@@ III Static Correlations
We start by determining the ground state spin configuration of FePSe3 by using the (adequately classically renormalized) exchanges listed in Table 1. We do so in an unbaised manner by performing Langevin dynamics simulations using the Sunny.jl package [48, 49, 50, 51] [See Appendix C for details]. Starting from the high temperature (random) limit, we thermalize a $24\times 24\times 2$ supercell of spins (in the hexagonal setting) followed by gradient descent to find the $T=0$ magnetic structure. We verify that for both FePSe3 models (BL and BQ), the ground state magnetic order is the $\boldsymbol{k}=(\frac{1}{2},0,\frac{1}{2})$ zig-zag structure, or its domains generated under $120^{\circ}$ rotations. In the case of FePS3, we find the $\boldsymbol{k}=(0,1,\frac{1}{2})$ structure, which is the same zig-zag arrangement as observed in FePSe3 up to a redefinition of the basis vectors as appropriate for the $C2/m$ monoclinic space group.
@@
To further explore the nature of the magnetic ordering transition in FePSe3 we calculate the magnetic order parameter as a function of temperature using the same Langevin dynamics approach. Care was taken to ensure samples were taken after sufficient time to thermalize and decorrelate the spin configurations [Appendix  D]. We observe a sharp transition at $T_{N}$ followed by a gradual saturation as $T\to 0$ [Appendix E].
@@
A well-known deficiency of classical simulations is the failure to capture the first-order nature of the transition. This fact was pointed out in Ref. 52 and can be overcome by restricting the dimensionality of the fluctuations manifold by considering Ising spins (and thus taking $J_{i}^{\alpha\beta}\to 0$ for $\alpha,\beta\neq z$). This leads to a better accounting of the first-order nature of this transition. However, such a model does not accurately describe the dynamics in systems with manifestly Heisenberg spins and exchange. Based on the fitting to the low temperature neutron data alone, the presence of Ising-like exchange anisotropy cannot be ruled out entirely, and the inclusion of such anisotropies may lead to a better description of the ordering transition.
@@
Figure 1: (a) The magnetic order parameter corresponding to $\boldsymbol{k}=(\frac{1}{2},0,\frac{1}{2}$) for the biquadratic (BQ) model of FePSe3 simulated using Langevin dynamics on a (24,24,2) supercell. A sharp phase transition is observed at $T_{N}^{\mathrm{cl}}=70$ K. (b) The in-plane magnetic structure with up (down) spins in purple (yellow). (c) Definitions of the first three nearest neighbor couplings.
The overestimation of fluctuations close to the ordering temperature also manifests in an underestimation of $T_{\rm N}$. In our classical simulations, we find that $T^{\mathrm{cl}}_{N}=70$ K, whilst experimentally, the ordering temperature was determined to be $T^{\mathrm{exp}}_{\rm N}=110$ K. This underestimation can be understood in the context of Weiss mean field theory which predicts $T_{N}$ to differ by a factor of $(S+1)/S$ between the quantum and classical models. Notwithstanding these limitations, our classical simulations correctly capture the sharp order parameter increase between a paramagnetic state and a zig-zag antiferromagnetic state below $T_{\rm N}$. Another contributing factor to these discrepancies is the use of Gaussian white noise which does not account for the necessary crossover to quantum (Bose-Einstein) statistics to describe spin fluctuations of the model at low temperatures. The implementation of colored noise [53] has been shown to accurately capture the low temperature thermodynamic properties of magnetic systems [54], but, as far as we know, a practical implementation of such thermostat in classical dynamics simulations is still lacking. Nevertheless, traditional Landau-Lifshitz dynamics has been shown to qualitatively reproduce the experimentally measured finite temperature dynamics of interacting spin systems with an appropriate rescaling of the temperature [see Refs. [55, 52, 56] for example].
@@ IV Linear Spin Wave Spectrum
A complete understanding of the spin dynamics of FePSe3 requires modeling dispersive higher energy spin-orbital excitations observed around $\sim$ 39 and 44 meV. In principle, this can be achieved by representing the local Hilbert space using coherent states of SU$(5)$ and including bilinear exchange interactions that directly couple components of the resulting magnetic multipoles. However, the coupling between high- and low-energy excitations is small, leading to a negligible neutron scattering intensity for these modes. Including mixing between higher spin-orbital states could potentially give rise to non-zero matrix elements between the ground-state and higher energy excited states; neutron scattering measurements at higher energy transfer may shed light on this question. Given the number of possible exchange interactions between SU(5) coherent states, we instead rely on a purely dipolar model for the spins in FePSe3.
@@
Table 1: Parameters used in our calculations, adapted using the RCS to be consistent with Chen et al. [1]. A small interchain coupling has been added to reproduce the reported magnetic ground state. ∗∗Reported value includes the RCS renormalization.
Bond BL BQ
$J_{1}$2.30 -1.32
$J_{2}$0.23  0.12
$J_{3}$  2.01  1.28
$J_{c}$  0.01  0.01
$K_{1}$  0.00 -0.22∗∗
$K_{2}$  0.00  0.00
$D$ 3.65  3.08
A cruder but more practically feasible approach is to introduce biquadratic exchange between purely dipolar spins, as it has been suggested to provide a better description of experimental data in FePSe3 [1], and similar arguments have been made for FePS3 [31]. We therefore restrict our further discussion to the BQ model. The inclusion of biquadratic coupling in $2D$ magnetism has been discussed at length in the literature (see, for example, Refs. [57, 58, 59, 60, 61]). Presence of this term does not protect order in $2D$ [62] but is often required for a strong agreement with experiment, particularly in the extreme two-dimensional limit [63, 64].
@@ V Temperature-dependent spin fluctuations
Figure 3: (a-g) Simulated inelastic neutron scattering intensity represented as $(\mathbf{Q},E)$ slices taken at increasing temperatures. Slices in the second row are plotted on a smaller intensity scale to aid visibility. Gray vertical lines indicate the location of the M points in the first Brillouin zone. Numerical divergences that appear as artifacts in the ordered phase due to the finite length of the real-time trajectories are deliberately obscured by these gray lines. (h) Constant energy cuts at $E=5$ meV taken through the $(\mathbf{Q},E)$ slice plotted in (a-g).
@@
Simulation results are plotted in Fig. 3. The calculations are once again domain-averaged, and are integrated over $L=[-2,2]$ out of plane. At $T=0.64T_{N}^{\mathrm{cl}}$, a dispersive mode is still visible, albeit with a broadened lineshape compared to the zero-temperature limit. The broad linewidth indicates a finite lifetime induced by thermal nonlinearities of the classical dynamics. At $T=0.86T_{N}^{\mathrm{cl}}$ the dispersive modes are still present with a reduction in intensity reflecting a redistribution of spectral weight to diffuse quasi-elastic features. A further reduction of the anisotropy gap is caused by a thermal fluctuation induced reduction of the ordered moment. As $T$ crosses the ordering temperature [Fig. 3(d)], broad features extend from the wavevectors associated with the antiferromagnetic ordering wavevector, indicating the presence of short-range antiferromagnetic correlations. These extend up and join the spin wave branch at $T_{N}$. The broad columnar features persist to higher temperatures [Figs. 3d-f] with the increased temperature redistributing intensity to the columns protruding from the locations of the magnetic Bragg peaks. The scattering remains structured at temperatures far exceeding the magnetic ordering temperature [Fig.3(f)], reproducing the signal observed experimentally [1]. To further clarify this temperature evolution, we calculate a constant energy cut at $E\!=\!5$ meV. At $T=0.64T_{N}^{\mathrm{cl}}$ no signal is observed in this range. As temperature is increased intensity accumulates at the antiferromagnetic wavevector leading to broad features across the Brilloin zone at $T=1.07T_{N}^{\mathrm{cl}}$, $T=1.08T_{N}^{\mathrm{cl}}$ and $T=1.14T_{N}^{\mathrm{cl}}$.
Present in our calculation is a region of high intensity at the zone center above $T_{N}$ which can be attributed to incipient short range spin wave excitations which develop close to $T_{N}$ and form the bottom of the magnon band seen in the ordered phase, where the density of thermally excited magnons is greatest [Fig. 3(a)]. In our calculation this feature is smeared out at higher temperatures. The evidence for this excess of intensity at the zone center in the experimental data [65] is inconclusive.